# Object Proxy Patterns for Accelerating Distributed Applications

### v1

@@ V-B Scalable Stream Processing
Figure 6: Compute tasks completed per second as a function of stream item data size and number of workers. One worker generates data consumed by a central dispatcher that launches simulated compute tasks (one second sleep tasks) for each item across the remaining $n-1$ workers. At small data sizes ($\leq 100$ KB), data transfer overheads are negligible and the dispatcher can keep up with incoming stream data; however, at large data sizes and worker counts, the dispatcher becomes overwhelmed by the size of data transfers required for each task. ProxyStream transparently decouples data flow from control flow improving overall system performance as stream data sizes and the number of workers is increased.
@@
Setup: We compare two streaming configurations. In Redis Pub/Sub, data are published directly to a Redis pub/sub topic that is consumed by the dispatcher before being sent to a worker to be computed on. In ProxyStream, data are published to a StreamProducer which decouples metadata from bulk data, sending metadata to a Redis pub/sub topic and storing bulk data in a Redis key/value store. The dispatcher consumes proxies of stream data via the StreamConsumer and sends proxies to workers to be computed on, thus avoiding expensive data transfers through the dispatcher.
@@
Results: Fig 6 shows the average compute tasks completed per second. For larger worker counts $n$ and data sizes $d$, the default Redis Pub/Sub deployment slows because the dispatcher becomes a bottleneck, processing stream data at 59 MB/s. This rate is slower than the network connection between the Redis server and dispatcher because the dispatcher must, for each stream item, receive and deserialize the item from Redis; compose the task payload, serializing the item again; and communicate the task payload to a worker. Thus the dispatcher cannot process the incoming stream data fast enough to keep workers fed with new tasks when the number of workers or data size is sufficiently high.
In contrast, ProxyStream transparently alleviates data transfer and serialization burdens from the dispatcher enabling better performance at scale. Stream data are only resolved on the compute node that executes the corresponding task, improving the latency between the dispatcher receiving stream data and launching a new task. We observe a steam data processed throughput of 1.1 GB/s with 1 MB items and 2048 workers and 1.2 GB/s with 100 MB items and 256 workers which represents a 20$\times$ increase in task throughput. The increase in performance with larger items and fewer workers is attributed to improved network utilization when sending few large messages as opposed to many small messages. We also note that at 1.2 GB/s, storing data produced by the generator is the limiting factor. A faster data storage system or multiple data generators would be needed to achieve scaling beyond this point, and ProxyStream does support modular data storage and multi-producer configurations.
@@ V-C Memory Management
Setup: We execute a simulated workflow that imitates a series of map-reduces across a local Dask cluster on a single compute node of Polaris. We run the workflow using each of the proxy memory management models, default, manual, and ownership, and a baseline without proxies using Dask for all data management. We record average memory usage across three workflow executions for each configuration. Eight consecutive map-reduces are performed where each of 32 mappers receives 100 MB and produces 10 MB of data. We choose 100 MB because the value is large enough to be observable in the memory trace (i.e., larger than the baseline memory usage fluctuations) but is also below the Redis default maximum value size of 512 MB. A single reducer consumes data produced by all mappers. In addition to consuming and producing data, each tasks sleeps for 5 s.
@@ VI Applications
We use the 1000 Genomes workflow to investigate ProxyFutures as a mechanism for reducing task overheads and extending data flow dependencies to FaaS systems. Each task in the original workflow is an executable program that reads and writes one or more intermediate files. We mitigate overheads associated with file system accesses by modifying the code to invoke functions instead, yielding a version that can be executed using a FaaS execution engine such as Globus Compute (which we use in the experiments reported here, due to its integration with HPC systems).
@@ VI-B DeepDriveMD
We modify the Parsl implementation of DeepDriveMD [50] to stream inference batches and results to and from a single, persistent inference task with ProxyStream. A persistent inference task eliminates task overheads and enables reuse of models and caches. Streaming with proxies reduces overheads in the DeepDriveMD client because received inference results are immediately added to a queue of simulation task inputs. In addition to ProxyStream, ProxyFutures are used to indicate availability of a new ML model to the inference task and proxy references for management of intermediate task data.
@@ VI-C MOF Generation
We modify the MOF Generation application to communicate all task input and output data larger than 1 kB via proxies. Note that the overhead of proxying simple data types such as boolean flags or configuration strings is greater than sending those objects directly. We deploy the application with default settings on ten Polaris nodes. We run the application twice: with the standard proxy implementation of ProxyStore and with our proxy ownership model. Here, ownership was sufficient; we did not use the lifetimes model. We record the number of actively proxied objects during the application’s runtime. As shown in Fig 10, the ownership model appropriately evicts proxied data when the lifetime of the associated proxy ends without altering the runtime behaviour of the application. Manual memory management is possible, as discussed in Sec V-C, but automated management is both safer and makes the adoption of advanced programming practices, such as those we present here, easier and more appealing.
@@ VII Related Work
In contrast, high-performance science applications often produce large raw or unstructured data accompanied by structured metadata [63]. File-oriented distributed applications often use GridFTP [41, 64]. Dispel4py [65, 66] maps abstract definitions of streaming workflows onto concrete distributed execution frameworks, such as Python multiprocessing or MPI [67]. Streamflow extends the DAG-based workflow model to integrate continuous event processing [68]. ADIOS WASP, a data staging platform for scientific stream processing, uses a self-describing file format and supports advanced networking technologies such as RDMA [69]. The SciStream middleware enables fast, secure memory-to-memory streaming between nodes that lack direct network connectivity [70].

### v2

@@ V-B Scalable Stream Processing
Figure 6: Compute tasks completed per second as a function of stream item data size and number of workers. One worker generates data consumed by a central dispatcher that launches simulated compute tasks (one second sleep tasks) for each item across the remaining $n-1$ workers. At small data sizes ($\leq 100$ KB), data transfer overheads are negligible and the dispatcher can keep up with incoming stream data; however, at large data sizes and worker counts, the dispatcher becomes overwhelmed by the size of data transfers required for each task in the Redis Pub/Sub configuration. ProxyStream transparently decouples data flow from control flow improving overall system performance as stream data sizes and the number of workers is increased.
@@
Setup: We compare three streaming configurations. In Redis Pub/Sub, data are published directly to a Redis pub/sub topic that is consumed by the dispatcher before being sent to a worker to be computed on. In ADIOS2, data are written step-by-step to an ADIOS2 stream [51]. The dispatcher iterates on steps and launches worker tasks which will read the data from the ADIOS2 stream at a specified step. In ProxyStream, data are published to a StreamProducer which decouples metadata from bulk data, sending metadata to a Redis Pub/Sub topic and storing bulk data in a Redis Key/Value store. The dispatcher consumes proxies of stream data via the StreamConsumer and sends proxies to workers to be computed on. ADIOS2 and ProxyStream avoid data transfers through the dispatcher.
@@
Results: Fig 6 shows the average compute tasks completed per second. At the smallest data size, $d=100$ kB, performance is comparable between the three methods because data are not large enough to stress the system. For larger worker counts $n$ and data sizes $d$, the default Redis Pub/Sub deployment slows because the dispatcher becomes a bottleneck, processing stream data at $\sim$100 MB/s. This rate is slower than the network connection between the Redis server and dispatcher because the dispatcher must, for each stream item, receive and deserialize the item from Redis; compose the task payload, serializing the item again; and communicate the task payload to a worker. Thus the dispatcher cannot process the incoming stream data fast enough to keep workers fed with new tasks when the number of workers or data size is sufficiently high.
ADIOS2 performs better than Redis Pub/Sub because we configured workers to read items from the stream directly based on a step index provided by the dispatcher, improving the latency between the dispatcher receiving stream data and launching a new task. However, ADIOS2 requires changes to the worker task code not needed by the other two methods.
ProxyStream also alleviates data transfer and serialization burdens from the dispatcher enabling performance on par with or better than ADIOS2 but does so transparently without needing changes to the worker task code. The peak processing throughput of ProxyStream is 1.7×1.7$\times$1.7 × and 2.0×2.0$\times$2.0 × faster than ADIOS2 for 1 MB and 10 MB item sizes, respectively. Compared to the Redis Pub/Sub baseline, ProxyStream is 4.6×4.6$\times$4.6 × and 6.2×6.2$\times$6.2 × faster for 1 MB and 10 MB item sizes, respectively. At $d=100$ MB, the largest data size evaluated, and $n=256$, ProxyStream is 7.3×7.3$\times$7.3 × faster than Redis Pub/Sub. ProxyStream and ADIOS2 perform similarly at this scale because other aspects of the experimental configuration become bottlenecks. Namely, task execution overheads and storing the data produced by the generator limit peak throughput. A faster data storage system or multiple data generators would be needed to achieve scaling beyond this point, and ProxyStream does support modular data storage and multi-producer configurations.
@@ V-C Memory Management
Setup: We execute a simulated workflow that imitates a series of map-reduces across a local Dask cluster on a single compute node of Polaris. We run the workflow using each of the proxy memory management models, default, manual, and ownership, and a baseline without proxies using Dask for all data management. We record average memory usage across three workflow executions for each configuration. Eight consecutive map-reduces are performed where each of 32 mappers receives 100 MB and produces 10 MB. We choose 100 MB because the value is large enough to be observable in the memory trace (i.e., larger than the baseline memory usage fluctuations) but is also below the Redis default maximum value size of 512 MB. A single reducer consumes data produced by all mappers. In addition to consuming and producing data, each tasks sleeps for 5 s.
@@ VI Applications
1000 Genomes: We use the 1000 Genomes workflow to investigate ProxyFutures as a mechanism for reducing task overheads and extending data flow dependencies to FaaS systems. Tasks in the original 1000 Genomes workflow were implemented as Bash scripts. We use the Python implementation of 1000 Genomes, where tasks are implemented as functions, to execute the workflow using a FaaS execution engine such as Globus Compute (which we use in the experiments reported here, due to its integration with HPC systems).
@@
DeepDriveMD: We modify the Parsl implementation of DeepDriveMD [53] to stream inference batches and results to and from a single, persistent inference task with ProxyStream. A persistent inference task eliminates task overheads and enables reuse of models and caches. Streaming with proxies reduces overheads in the DeepDriveMD client because received inference results are immediately added to a queue of simulation task inputs. In addition to ProxyStream, ProxyFutures are used to indicate availability of a new ML model to the inference task and proxy references for management of intermediate task data.
@@
MOF Generation: We modify the MOF Generation application to communicate all task input and output data larger than 1 kB via proxies. (The overhead of proxying simple data types such as boolean flags or configuration strings is greater than sending those objects directly.) We deploy the application with default settings on ten Polaris nodes. We run the application twice: with the standard proxy implementation of ProxyStore and with our proxy ownership model. Here, ownership was sufficient; we did not use the lifetimes model. We record the number of actively proxied objects during the application’s runtime. As shown in Fig 10, the ownership model appropriately evicts proxied data when the lifetime of the associated proxy ends without altering the runtime behaviour of the application. Manual memory management is possible, as discussed in Sec V-C, but automated management is safer and makes adoption of advanced programming practices, such as those we present here, easier and more appealing.
@@ VII Related Work
In contrast, high-performance science applications often produce large raw or unstructured data accompanied by structured metadata [65]. File-oriented distributed applications often use GridFTP [39, 66]. Dispel4py [67, 68] maps abstract definitions of streaming workflows onto concrete distributed execution frameworks, such as Python multiprocessing or MPI [69]. Streamflow [70] extends the DAG-based workflow model to integrate continuous event processing. ADIOS WASP [71], a data staging platform for scientific stream processing, uses a self-describing file format and supports advanced networking technologies such as RDMA. The SciStream middleware [72] enables fast, secure memory-to-memory streaming between nodes that lack direct network connectivity. CAPIO [73] provides a middleware layer for injecting I/O streaming capabilities into file-based workflows.